The space shuttle endeavor has a mass of 2.0x10^6 kg as it sit on the launching pad at cape canaveral The rocket engines have a combined thrust of 3.0x10^7 at take off.

A)What is the weight of the rocket on the launch pad?
B)What is the acceleration of the Endeavor as it takes off?
C)The average acceleration of the Endeavor during it's 10 minute launch is 13m/s^2. What is the final velocity during this time period?
D)As the shuttle flies higher and higher, would you expect the acceleration to decrease, stay the same, or increase? The thrust from the engines remains the same during the flight. Explain!
1
Expert's answer
2019-03-15T12:45:06-0400

(A) The weight of rocket

"W=mg=2.0\times 10^6\:\rm{kg}\times 9.8\:\rm{N/kg}=1.96\times 10^7\:\rm{N}"

(B) The acceleration of the Endeavor


"a=\frac{F_{\rm{thrust}}-W}{m}=\frac{3.0\times 10^7\:\rm{N}-1.96\times 10^7\;\rm{N}}{2.0\times 10^6\:\rm{kg}}=5.2\:\rm{m/s^2}"

(C) The final velocity

"v_f=v_i+at=0\:\rm{m/s}+13\:\rm{m/s^2}\times (10\times 60\:\rm{s})=7800 \:\rm{m/s}"

(D) Since the acceleration due gravity decreases with height, so the acceleration of the Endeavor will increase.